# Step-DeepResearch Technical Report

### v1

@@ 7.2 Evaluation Results
Figure 6 presents a comprehensive comparison between Step-DeepResearch and other agents on Research Rubrics benchmark. The experimental results demonstrate that Step-DeepResearch achieves SOTA performance within the single-agent category. With a score of $61.42$, it significantly outperforms the leading open-source model, Kimi-k2-thinking ($56.17$), representing a $5.25$ improvement in performance. In the overall leaderboard, Step-DeepResearch ranks second only to the commercial system Gemini DeepResearch $63.69$, surpassing most complex agent frameworks including OpenAI DeepResearch and Kimi-Researcher. Notably, the superior performance of Step-DeepResearch is characterized by exceptional parameter efficiency:
@@
  • Lightweight Advantage: Among commercial agent systems, high-end models incur significant costs per report: Gemini DeepResearch ($\approx$ 6.65 RMB) and OpenAI DeepResearch ($\approx$ 5.32 RMB) are the most expensive, followed by MiniMax Agent Pro ($\approx$ 3.36 RMB) and Kimi-Researcher ($\approx$ 2.66 RMB), with Qwen DeepResearch being the outlier ($\approx$ 0.63 RMB). In the ReAct agent category, costs are generally lower but variable: MiniMax-M2 ($\approx$ 0.42 RMB), DeepSeek-V3.2 ($\approx$ 0.68 RMB), Kimi-k2-thinking ($\approx$ 0.76 RMB), and GLM-4.6 ($\approx$ 1.05 RMB). Remarkably, Step-DeepResearch achieves a ResearchRubrics score of 61.42 with a single invocation cost of less than 0.50 RMB. This represents a cost reduction of more than 10-fold compared to top-tier commercial systems (e.g., Gemini and OpenAI) while maintaining state-of-the-art performance, demonstrating exceptional cost-effectiveness for large-scale deployment. Cost metrics are shown in Figure 1(a).
@@
ADR-Bench.
As shown in Figure 7, Step-DeepResearch demonstrates superior performance in human evaluations compared to existing commercial agent systems. In an ablation study comparing Step-DeepResearch with its non-midtrained version, the model achieved a record of $30$ wins and $21$ losses. This result indicates that the integration of mid-training specifically for agents aligns more closely with human preferences regarding the quality of complex research reports. When benchmarked against other leading systems, Step-DeepResearch consistently maintains a higher win rate than loss rate, establishing its position as a superior alternative to current systems. Notably, the model demonstrates significant non-inferiority across all benchmarks. Specifically, against formidable opponents such as Gemini and MiniMax, the cumulative count of “Wins” and “Ties” reached $47$ ($67.1\%$). These findings provide empirical evidence that Step-DeepResearch consistently meets or exceeds the most advanced performance standards across the vast majority of research scenarios.
@@
Queries from finance and law are typically feature industry terminology, multi-stage reasoning chains, and sensitive constraints. To address this, we incorporated explicit negative scoring criteria within the checklist for each data point to penalize errors in reports. For unpardonable errors, the penalty is particularly severe, leading to a direct “unusable” (zero-score) rating for that specific report. Under this scoring framework, a model’s overall performance is no longer driven solely by its hit rate but also reflects the cost of errors and their associated risk. Evaluation results shown in table 6, reveal a clear tri-tier distribution of performance. Gemini outperformed all others, establishing a dominant lead, Step-DeepResearch, Kimi, and OpenAI showed comparable performance, clustering within the same tier. The remaining models and products lagged significantly behind, exhibiting a pronounced long-tail effect. These findings suggest that scoring variances are more likely correlated with the models’ inherent domain-knowledge coverage rather than gains from agentic frameworks. In other words, under the constraint of strict negative scoring, the process optimizations provided by Agent frameworks cannot compensate for a model’s fundamental knowledge gaps. Step-DeepResearch’s position at the forefront of the second tier is precisely due to its domain-specific training in financial and legal scenarios, granting it a level of expertise competitive with much larger parameter models.
@@ 7.3 Detailed Analysis
Figure 8. shows the performance distribution of Step-DeepResearch (denoted by blue diagonal-hatched bars) across six critical evaluation dimensions. The experimental results highlight several key performance breakthroughs: Step-DeepResearch demonstrated significant leadership in “Implicit Criteria” and “Explicit Criteria” securing scores of 54.5 and 72.0, respectively. Notably, its performance in implicit criteria surpasses OpenAI DeepResearch (52.4) and substantially exceeds open-source models, signaling a robust capacity for underlying logical alignment. In the category of “Citation Quality” Step-DeepResearch achieved a score of 57.0, tying for the top position with Gemini DeepResearch. This high level of empirical rigor ensures that all generated insights are fully substantiated by verifiable sources. With a score of 58.2 in “Communication Quality”, the model outperformed all evaluated counterparts. This indicates that the generated reports are not only data-rich but also exhibit superior clarity and professional readability. Comparative analysis also identified a marginal gap in “Instruction Following”, where Step-DeepResearch (64.9) remains slightly behind Kimi-Researcher (66.7). Our preliminary diagnostic attributes this to the breadth of instruction fine-tuning data diversity during the Post-training phase. Future development will prioritize targeted optimization for multi-constraint complex tasks to ensure a dominant lead across all evaluative dimensions. As shown in Figure 9, Step-DeepResearch outperformed Kimi-Researcher in AI & ML (64.8 vs 57.4), Historical Analysis (65.8 vs 61.4), and Technical Documentation (64.6 vs 53.1), tying with Gemini DeepResearch for the top spot in the latter. Notably, these results were achieved without domain-specific data augmentation. The model also led in Creative Writing (63.4), demonstrating that its underlying logic and knowledge structure generalize effectively to out-of-distribution tasks. Despite broad leadership, the model shows room for improvement in STEM (64.7) and Philosophy (46.1), currently trailing Gemini DeepResearch (70.7 and 53.5, respectively). Error analysis indicates these gaps stem from the high demands of high-order reasoning. Future efforts will focus on closing these logical reasoning gaps to enhance performance in hard-science domains.
@@
We employed a fine-grained, pairwise human preference evaluation methodology to conduct a comprehensive performance alignment test of Step-DeepResearch, shown in Figure.10. A detailed analysis of each dimension follows below:
  • Informational Completeness. For tasks involving cross-domain knowledge synthesis, evidence chain tracking, and multi-source information aggregation, the model not only covers all key points but also provides structured expression, enhancing both information capacity and perceived quality. In tasks requiring large-scale categorization or element listing, the model significantly suppresses content hallucinations and redundant stacking achieving a complete and accurate structured presentation.
  • Content Depth. Regarding content depth, Step-DeepResearch achieves performance comparable to several closed-source commercial systems. In scenarios that heavily rely on specialized professional knowledge, it trails slightly behind DeepResearch-style products such as OpenAI and Gemini, a gap we attribute primarily to differences in model parameter scale. By contrast, Step-DeepResearch demonstrates more consistent execution capabilities in general problem analysis, planning, and strategic inference tasks. We observe that base models without task-specific training often exhibit shallow response patterns, characterized by short, loosely connected sentences and superficial bullet points that provide limited practical value. Step-DeepResearch effectively mitigates this behavior through targeted training.
  • Requirement Fitness. In terms of meeting user requirements, Step-DeepResearch demonstrates a level of stability comparable to that of commercial systems. Across a wide range of instruction complexities, the model is able to accurately identify user intent and provide relevant supplementary information to support decision-making. We observe that several existing DeepResearch-style systems rely on rigid output patterns, such as enforcing a thesis-like structure across diverse tasks or introducing information that is only weakly related to the user’s intent. In contrast, Step-DeepResearch adapts its output structure to the task at hand, allowing writing style and presentation to evolve dynamically based on task attributes.
@@ 7.4 Case study
Table 7 presents a representative example from our benchmark comparing midtrained models with non-midtrained baselines. The results demonstrate substantial improvements in overall report quality for midtrained models. In particular, we observe consistent gains in instruction following, as well as in the breadth and depth of information exploration. These improvements are strongly correlated with the incorporation of atomic capabilities, such as planning and information seeking during the midtraining phase.
@@
English Translation. Regarding repo-level code engineering capabilities of large language models (similar to the SWE-bench evaluation), many studies propose leveraging GitHub PRs and Issues to construct midtraining corpora between pre-training and post-training. While SWE-RL provides a detailed data construction and rewriting pipeline, I seek other works that (1) explicitly describe PR/Issue data construction, (2) release open-source data cleaning and reconstruction code, or (3) provide reconstructed datasets. I am particularly interested in filtering criteria, cleaning procedures, and format conversion methods. Please systematically review related works (e.g., SWE-Fixer, SWE-agent, DeepSWE), with an emphasis on those that offer detailed schemes for organizing and rewriting PR, Issue, and repo-level data.
@@
Information Completeness Covers over a dozen relevant works, including the SWE-bench series, SWE-smith, SWE-Factory, Multi-SWE-bench, SWE-RL, DeepSWE, InstructCoder, Coeditor, RepoBench, SWE-agent-trajectories, and SWE-Fixer. Covers only six works, omitting foundational datasets and key automation frameworks such as SWE-smith and SWE-Factory.
Content Depth Provides detailed, end-to-end data processing pipelines, concrete code-level examples, and in-depth discussions of key mechanisms such as Issue-PR pairing, test-case isolation, and automated environment construction. Presents structured comparative tables summarizing data sources, formats, filtering criteria, cleaning steps, format conversions, and open-source availability. Offers deeper analysis for a small subset of works (e.g., SWE-RL) but discusses most studies only at a high level. Lacks systematic cross-paper comparisons; information is presented in a fragmented, narrative-driven manner.
@@ 7.5 Targeted Refinements for Deep Research Patterns
While earlier models achieved high Research Rubric scores ($57$), human evaluation revealed a tendency toward fragmented information stacking rather than deep analysis. This highlights a limitation in checklist-based scoring, which prioritizes information stacking rather than deep, integrated analysis. This discrepancy highlights a limitation in checklist-based metrics: they prioritize information recall (i.e., whether key points exist) over structural coherence and narrative depth. Furthermore, we identified a negative correlation between analytical depth and comprehensiveness. When prompted to minimize list-based formatting and provide deeper insights, the model often terminated generation prematurely, leading to a drop in recall. To address this, we optimized our data synthesis pipeline by introducing a Synthesis-driven Drafting module. This module directs the model to transform raw tool-call trajectories into structured paragraphs with logical deduction, while strictly limiting the use of unordered lists as the primary content body. Additionally, we implemented a Pairwise LLM Judger for quality control. This filtering mechanism retains new reports only if they demonstrate superior depth and logic without compromising overall information coverage.
@@
To address inconsistent Chinese-English code-switching that disrupts readability, we applied a a strict data-cleaning strategy. During the trajectory synthesis phase, we utilize a combination of regular expressions and language density detection to identify and remove low-quality trajectories. This process programmatically filters out instances of non-essential language mixing—specifically the interspersing of common English vocabulary within Chinese prose—while preserving necessary proper nouns and domain-specific technical terminology.
@@ 8 Conclusion and Future Work
To address these issues, our future work will advance along three primary directions. First, we will develop a multi-agent collaborative paradigm by introducing specialized roles such as planners, retrievers, verifiers, and writers to reduce hallucinations through consensus mechanisms. Second, we will expand to complex environment interactions, enabling the agent to perform continuous exploration and error correction in partially observable and dynamic settings. Third, we will improve reward modeling and training objectives to incorporate metrics such as factual consistency, traceable citations, and structural clarity. By combining preference learning with process-based supervision, we aim to ensure that the model is not only “seemingly correct” but “verifiably and clearly correct.”
@@ Acknowledgements
Appendix A Appendix
@@
Category Chinese Prompt
Social Life Scenario: 多约束复杂旅行规划:带娃、多人、出国、城市移动旅行方案。

### v2

@@ 7.2 Evaluation Results
Figure 6 presents a comprehensive comparison between Step-DeepResearch and other agents on ResearchRubrics benchmark. The experimental results demonstrate that Step-DeepResearch achieves state-of-the-art performance within the single-agent category. With a score of $61.42$, it significantly outperforms the leading open-source model, Kimi-k2-thinking ($56.17$), representing a $5.25$ improvement in performance. In the overall leaderboard, Step-DeepResearch ranks second only to the commercial system Gemini DeepResearch $63.69$, surpassing most complex agent frameworks including OpenAI DeepResearch and Kimi-Researcher. Notably, the superior performance of Step-DeepResearch is characterized by exceptional parameter efficiency:
@@
  • Lightweight Advantage: Among commercial agent systems, high-end models incur significant costs per report: Gemini DeepResearch ($\approx$ 6.65 RMB) and OpenAI DeepResearch ($\approx$ 5.32 RMB) are the most expensive, followed by MiniMax Agent Pro ($\approx$ 3.36 RMB) and Kimi-Researcher ($\approx$ 2.66 RMB), with Qwen DeepResearch being the outlier ($\approx$ 0.63 RMB). In the ReAct agent category, costs are generally lower but variable: MiniMax-M2 ($\approx$ 0.42 RMB), DeepSeek-V3.2 ($\approx$ 0.68 RMB), Kimi-k2-thinking ($\approx$ 0.76 RMB), and GLM-4.6 ($\approx$ 1.05 RMB). Remarkably, Step-DeepResearch achieves a ResearchRubrics score of 61.42 with a single invocation cost of less than 0.50 RMB. At less than one-tenth the expense of top-tier commercial systems such as Gemini and OpenAI, it maintains state-of-the-art performance, demonstrating exceptional cost-effectiveness for large-scale deployment. Cost metrics are shown in Figure 1(a).
ADR-Bench.
As shown in Figure 7, Step-DeepResearch demonstrates superior performance in human evaluations compared to existing commercial agent systems. In an ablation study comparing Step-DeepResearch with its non-midtrained version, the model achieved a record of $30$ wins and $21$ losses. This result indicates that the integration of mid-training specifically for agents aligns more closely with human preferences regarding the quality of complex research reports.
When benchmarked against other leading systems, Step-DeepResearch consistently maintains a higher win rate than loss rate, establishing its position as a superior alternative to current systems. Notably, the model demonstrates significant non-inferiority across all comparisons. Specifically, against formidable opponents such as Gemini and MiniMax, the cumulative count of “Wins” and “Ties” reached $47$ ($67.1\%$). These findings provide empirical evidence that Step-DeepResearch consistently meets or exceeds the most advanced performance standards across the vast majority of research scenarios.
@@
Specifically for the finance and law subsets within ADR-Bench, the tasks are characterized by high-density industry terminology, multi-stage reasoning chains, and strict compliance risk constraints. To address these complexities, we incorporated explicit negative scoring criteria in our checklists to penalize any professional misconceptions identified in the generated reports. For errors deemed fatal or professionally critical by experts, the penalty is particularly severe, leading to a direct “unusable” (zero-score) rating for that specific report. Under this scoring framework, a model’s overall performance is no longer driven solely by its hit rate but also reflects the cost of errors and the externalities of risk. Evaluation results shown in table 6, reveal a clear tri-tier distribution of performance. Gemini outperformed all others, establishing a dominant lead, Step-DeepResearch, Kimi, and OpenAI showed comparable performance, clustering within the same tier. The remaining models and products lagged significantly behind, exhibiting a pronounced long-tail effect. These findings suggest that scoring variances are more likely correlated with the models’ inherent domain-knowledge coverage rather than gains from agentic frameworks. In other words, under the constraint of strict negative scoring, the process optimizations provided by agent frameworks cannot compensate for a model’s fundamental knowledge gaps. Step-DeepResearch’s position at the forefront of the second tier is precisely due to its domain-specific training in financial and legal scenarios, granting it a level of expertise competitive with much larger parameter models.
@@ 7.3 Detailed Analysis
Figure 8. shows the performance distribution of Step-DeepResearch (denoted by blue diagonal-hatched bars) across six critical evaluation dimensions. The experimental results highlight several key performance breakthroughs: Step-DeepResearch demonstrated significant leadership in “Implicit Criteria” and “Explicit Criteria” securing scores of 54.5 and 72.0, respectively. Notably, its performance in implicit criteria surpasses OpenAI DeepResearch (52.4) and substantially exceeds open-source models, signaling a robust capacity for underlying logical alignment.
@@
In the category of “Citation Quality” Step-DeepResearch achieved a score of 57.0, tying for the top position with Gemini DeepResearch. This high level of empirical rigor ensures that all generated insights are fully substantiated by verifiable sources. With a score of 58.2 in “Communication Quality”, the model outperformed all evaluated counterparts. This indicates that the generated reports are not only data-rich but also exhibit superior clarity and professional readability. Comparative analysis also identified a marginal gap in “Instruction Following”, where Step-DeepResearch (64.9) remains slightly behind Kimi-Researcher (66.7). Our preliminary analysis attributes this to the extensive diversity of instruction fine-tuning data during the Post-training phase. Moving forward, we will focus on specialized optimizations for multi-constraint complex tasks to achieve a comprehensive lead across all evaluative dimensions.
@@
As shown in Figure 9, Step-DeepResearch outperformed Kimi-Researcher in AI & ML (64.8 vs 57.4), Historical Analysis (65.8 vs 61.4), and Technical Documentation (64.6 vs 53.1), essentially tying with Gemini DeepResearch for the top spot in the latter. Notably, these results were achieved without domain-specific data augmentation. The model also led in Creative Writing (63.4), demonstrating that its underlying logic and knowledge structure generalize effectively to out-of-distribution tasks. Despite broad leadership, the model shows room for improvement in STEM (64.7) and Philosophy (46.1), currently trailing Gemini DeepResearch (70.7 and 53.5, respectively). Error analysis indicates these gaps stem from the inherent complexity of high-order reasoning. Future efforts will prioritize bridging these logical bottlenecks to enhance model performance across STEM and other highly technical domains.
To assess Step-DeepResearch in general domains, we employed a fine-grained, pairwise human preference evaluation methodology to conduct a comprehensive performance alignment test on Step-DeepResearch, shown in Figure.10. The following sections present a granular analysis of each performance dimension:
  • Informational Completeness. For tasks involving cross-domain knowledge synthesis, evidence chain tracking, and multi-source information aggregation, the model not only covers all key points but also provides a structured presentation, enhancing both information capacity and perceived quality. In tasks requiring large-scale categorization or element listing, the model significantly suppresses content hallucinations and redundant stacking, thereby ensuring a comprehensive and precise output.
  • Content Depth. Regarding content depth, Step-DeepResearch achieves performance comparable to several closed-source commercial systems. In scenarios that heavily rely on specialized professional reasoning, it trails slightly behind commercial Deep Research agents such as OpenAI and Gemini, a performance gap we attribute primarily to constraints in model parameter scale. By contrast, Step-DeepResearch demonstrates more consistent execution capabilities in general problem analysis, planning, and strategic inference tasks. We observe that base models without task-specific training often exhibit shallow response patterns, characterized by short, loosely connected sentences and superficial bullet points that provide limited practical value. Step-DeepResearch effectively rectifies these deficiencies through targeted training.
  • Requirement Fitness. In terms of meeting user requirements, Step-DeepResearch demonstrates a level of stability comparable to that of commercial systems. Across a wide range of instruction complexities, the model is able to accurately identify user intent and provide relevant supplementary information to support decision-making. We observe that several existing commercial systems rely on rigid output patterns, such as imposing a rigid academic structure across diverse tasks or introducing information that is only weakly related to the user’s intent. In contrast, Step-DeepResearch dynamically adapts its presentation style to the specific attributes of the task, ensuring that the output structure remains highly relevant and purpose-driven.
@@ 7.4 Case study
Table 7 presents a representative example from our benchmark comparing the mid-trained model against its non-mid-trained version. The results demonstrate substantial improvements in overall report quality for the mid-trained model. In particular, we observe consistent gains in instruction following, as well as in the breadth and depth of information exploration. These improvements are strongly correlated with the incorporation of atomic capabilities, such as planning and information seeking during the mid-training phase.
@@
English Translation. Regarding the repo-level code engineering capabilities of large language models (similar to the SWE-bench evaluation), many studies have proposed leveraging GitHub PRs and Issues to construct mid-training corpora between pre-training and post-training phases. SWE-RL provides a detailed data construction/rewriting process and proposes the concept of "mid-training" (though this is questionable, as I am not certain if it was the first to introduce the concept of code mid-training). I am interested in knowing if there are other works that provide detailed construction processes for GitHub PR and issue data, open-source the code for cleaning and reconstructing such data, or provide open-source reconstructed datasets. I am specifically focused on data processing methods, including selection criteria, cleaning steps, and format conversion for PR/Issue data. Please introduce all relevant works (e.g., SWE-Fixer, SWE-agent, DeepSWE), with an emphasis on those that offer detailed schemes for organizing and rewriting PR, Issue, and repo-level data.
@@
Information Completeness Covers a broader range of relevant works, including the SWE-bench series, SWE-smith, SWE-Factory, Multi-SWE-bench, SWE-RL, DeepSWE, InstructCoder, Coeditor, RepoBench, SWE-agent-trajectories, and SWE-Fixer. Covers only six works, omitting foundational datasets and key automation frameworks such as SWE-smith and SWE-Factory.
Content Depth Provides detailed, end-to-end data processing pipelines, concrete code-level examples, and in-depth discussions of key mechanisms such as Issue-PR pairing, test-case isolation, and automated environment construction. Presents structured comparative tables summarizing data sources, formats, filtering criteria, cleaning steps, format conversions, and open-source availability. While providing asymmetrical depth on a select few works (e.g., SWE-RL), the response offers only superficial coverage of the remaining studies. It lacks systematic cross-paper synthesis, presenting information in a fragmented, narrative-driven manner that fails to highlight technical trade-offs.
@@ 7.5 Targeted Refinements for Deep Research Patterns
While earlier models achieved high ResearchRubrics scores ($57$), human evaluation revealed a tendency toward fragmented information stacking rather than deep analysis. This highlights a limitation in checklist-based scoring, which prioritizes information stacking rather than deep, integrated analysis. This discrepancy highlights a limitation in checklist-based metrics: they prioritize information recall (i.e., whether key points exist) over structural coherence and narrative depth. Furthermore, we identified a negative correlation between analytical depth and comprehensiveness. When prompted to minimize list-based formatting and provide deeper insights, the model often terminated generation prematurely, leading to a drop in recall. To address this, we optimized our data synthesis pipeline by introducing a Synthesis-driven Drafting module. This module directs the model to transform raw tool-call trajectories into structured paragraphs with logical deduction, while strictly limiting the use of unordered lists as the primary content body. Additionally, we implemented a Pairwise LLM Judger for quality control. This filtering mechanism retains new reports only if they demonstrate superior depth and logic without compromising overall information coverage.
@@
To address inconsistent Chinese-English code-switching that disrupts readability, we applied a strict data-cleaning strategy. During the trajectory synthesis phase, we utilize regular expressions and language density detection to exclude low-quality trajectories characterized by unnecessary language mixing, ensuring a consistent and fluent output.
@@ 8 Conclusion and Future Work
Our future research aims to address these challenges through three strategic advancements. Initially, we focus on collaborative intelligence, introducing a multi-agent paradigm where specialized roles—planners, retrievers, verifiers, and writers—reduce hallucinations via consensus mechanisms. Beyond architectural design, we seek to enhance environmental adaptability, enabling agents to perform continuous exploration and error correction in dynamic, partially observable settings. Finally, we prioritize the rigor of training objectives by incorporating metrics such as factual consistency and traceable citations. Through the synergy of preference learning and process-based supervision, we strive to ensure that model outputs are not only “seemingly correct” but “verifiably and clearly correct.”
@@ Acknowledgements
Appendix
A. ResearchRubrics Judger System Prompt
@@ B. Examples
[…]